# AdaNCA: Neural Cellular Automata As Adaptors For More Robust Vision Transformer

### v1

@@ 3.1 Preliminaries
Vision Transformers (ViTs) operate on token maps $\mathbf{X}\in\mathbb{R}^{N\times C}$, where the number of tokens is $N$ and the each token is represented by a $C$-dimensional vector. ViTs learn the interaction between these tokens via self-attention [63] and computes the interaction result $\mathbf{X}_{attn}$, as described in Equation 1.
@@
$f$ is the MLP and $\theta$ stands for its parameters. The self-attention and MLP forms a single ViT block, and one can build a ViT model via stacking ViT blocks.
@@ 3.3 Insert positions of AdaNCA
Given a ViT and an AdaNCA, we need to determine where to insert AdaNCA to maximize the robustness improvements. To this end, we first establish the correlation between the placement of AdaNCA and the robustness enhancement it brings. Motivated by the fact that layer redundancy contributes to robustness [27], we hypothesize that the effect of AdaNCA should correlate to the layer redundancy corresponding to the insert position. To quantify the redundancy, we propose the Set Cohesion Index $\kappa$. Given a trained model with $L$ layers and two layer indices $i,j\in\{1,2,...,L\}$ where $i<j$, $\kappa(i,j)$ is defined in Equation 10.
@@ 7 Conclusion
We propose AdaNCA, an efficient Neural Cellular Automata (NCA) that, when inserted between their middle layers, improves ViT performances and robustness against adversarial attacks as well as out-of-distribution inputs. We motivate our design by connecting NCA and ViT in terms of token interaction modeling, and propose Dynamic Interaction to improve the computational efficiency of standard NCA. Exploiting the training strategies and design choices in NCA, i.e., stochastic update, random steps, and multi-scale interaction, we further enhance the AdaNCA-enhanced ViTs’ clean accuracy and robustness. To decide the placement of AdaNCA, we propose the Set Cohesion Index that quantifies the network redundancy via layer similarity and conclude that AdaNCA should be inserted between two layer sets that consist of redundant layers. Our results demonstrate that AdaNCA consistently improves ViTs performances and robustness. Evidence suggests that the mechanism by which we obtain improvement reduces the sensitivity of ViTs to certain types of noise and makes the noise-resilient ability of ViTs closer to humans.
@@ C.4 The effect of AdaNCA placement on robustness on ImageNet1K
Our algorithm for deciding the placement of AdaNCA is based on the correlation between the robustness improvement and the Stage Cohesion Index in Section 3.3. All the experiments are conducted on ImageNet100 to efficiently use the computational resources. Here, we compare two schemes for deciding the placement of AdaNCA on ImageNet1K on Swin-B [37] and FAN-B [77] to demonstrate our result validity. The first scheme is No-Prior, which does not involve the knowledge of the layer similarity and performs the most reasonable placement choice. For Swin-B, the choice for placing 3 AdaNCA is to insert them between each stage pair defined by the transition between different embedding dimensionalities, namely after layer 2, layer 4, and layer 22. Coincidentally, this choice aligns with what we obtain from the dynamic programming algorithm, which is to place AdaNCA after layer 2, layer 8, and layer 22. For FAN-B, although it is a hybrid ViT, we only consider inserting AdaNCA into its main network where all layers share the same structure. In this case, the No-Prior chooses to uniformly insert AdaNCA between the 16 layers, that is after layers 5 and 10. Our algorithm decides the placement should instead be after layers 6 and 9. Table 7 shows the results of two Swin-B-AdaNCA models. Inserting AdaNCA with the No-Prior scheme can still contribute to the model performance and robustness, but not as effective as using our proposed method.
@@ C.6 Training details
Table 8: Hyperparameters used in AdaNCA-equipped ViT training.
@@
Table 9: Training time of each AdaNCA-equipped ViT.
@@ C.7.2 Black-box attack
C.8 Reason for using ImageNet22K model for RVT
@@
C.9 Visualization of attention maps
@@
C.10 Additional results on same model architecture but with more layers ($\star$ models)
@@
C.11 Additional results on layer similarity structure
@@
Figure 10: Layer similarity structures of different ViTs. Stage marked in red boxes. AdaNCA is inserted between each pair of stages. $\kappa_{mean}$ is the mean of $\kappa$ of all stages marked out.
C.12 Additional results on category-wise mCE on ImageNet-C
@@
C.13 Additional results on noise sensitivity examination
@@
C.14 Datasets information and license
@@
C.15 Model and code license

### v4

@@ 3.1 Preliminaries
Vision Transformers (ViTs) operate on token maps $\mathbf{X}\in\mathbb{R}^{N\times C}$, where the number of tokens is $N$ and the each token is represented by a $C$-dimensional vector. ViTs learn the interaction between these tokens via self-attention [63] and computes the interaction result $\mathbf{X}_{attn}$, as described in Equation $1$We ignore the residual update scheme for simplicity.
@@
$f$ is the MLP and $\theta$ stands for its parameters. The self-attention and MLP form a single ViT block, and one can build a ViT model via stacking ViT blocks.
@@ 3.3 Insert positions of AdaNCA
Given a ViT and an AdaNCA, we need to determine where to insert AdaNCA to maximize the robustness improvements. To this end, we first establish the correlation between the placement of AdaNCA and the robustness enhancement it brings. Motivated by the fact that the network redundancy contributes to the model robustness [27], we hypothesize that the effect of AdaNCA should correlate to the layer redundancy corresponding to the insert position. To quantify the redundancy, we propose the Set Cohesion Index $\kappa$. Given a trained model with $L$ layers and two layer indices $i,j\in\{1,2,...,L\}$ where $i<j$, $\kappa(i,j)$ is defined in Equation 10.
@@ 7 Conclusion
We propose AdaNCA, an efficient Neural Cellular Automata (NCA) that, when inserted between their middle layers, improves ViT performances and robustness against adversarial attacks as well as out-of-distribution inputs. We motivate our design by connecting NCA and ViT in terms of token interaction modeling, and propose Dynamic Interaction to improve the computational efficiency of standard NCA. Exploiting the training strategies and design choices in NCA, i.e., stochastic update, random steps, and multi-scale interaction, we further improve the AdaNCA-enhanced ViTs’ clean accuracy and robustness. To decide the placement of AdaNCA, we propose the Set Cohesion Index that quantifies the network redundancy via layer similarity and conclude that AdaNCA should be inserted between two layer sets that consist of redundant layers. Our results demonstrate that AdaNCA consistently improves ViTs performances and robustness. Evidence suggests that the mechanism by which we obtain improvement reduces the sensitivity of ViTs to certain types of noise and makes the noise-resilient ability of ViTs closer to humans.
@@ C.4 The effect of AdaNCA placement on robustness on ImageNet1K
Our algorithm for deciding the placement of AdaNCA is based on the correlation between the robustness improvement and the Set Cohesion Index in Section 3.3. All the experiments are conducted on ImageNet100 to efficiently use the computational resources. Here, we compare two schemes for deciding the placement of AdaNCA on ImageNet1K on Swin-B [37] and FAN-B [77] to demonstrate our result validity. The first scheme is No-Prior, which does not involve the knowledge of the layer similarity and performs the most reasonable placement choice. For Swin-B, the choice for placing 3 AdaNCA is to insert them between each stage pair defined by the transition between different embedding dimensionalities, namely after layer 2, layer 4, and layer 22. Coincidentally, this choice aligns with what we obtain from the dynamic programming algorithm, which is to place AdaNCA after layer 2, layer 8, and layer 22. For FAN-B, although it is a hybrid ViT, we only consider inserting AdaNCA into its main network where all layers share the same structure. In this case, the No-Prior chooses to uniformly insert AdaNCA between the 16 layers, that is after layers 5 and 10. Our algorithm decides the placement should instead be after layers 6 and 9. Table 7 shows the results of two Swin-B-AdaNCA models. Inserting AdaNCA with the No-Prior scheme can still contribute to the model performance and robustness, but not as effective as using our proposed method.
@@ C.6 Training details
Table 8: Hyperparameters used in AdaNCA-equipped ViT training.
@@
Table 9: Training time of each AdaNCA-equipped ViT.
@@ C.7.2 Black-box attack
C.8 The effect of the range of the steps in random step training
Our random stap training strategy contributes to the model robustness as shown in Table 3. In all of our experiments, we adopt a range of 2 in the random step setting. Here, we examine the effect of increasing the range. We use the same setting as described in Section 4.3. Results are given in Table 12. We can observe that more choices of the recurrent steps worsen the model performance. This might stem from too much noise introduced into the training process which leads to underfitting. Therefore, we only adopt a range of 2 in our experiments.
Table 12: Results of Swin-Tiny and the AdaNCA-enhanced models with different ranges of random steps on ImageNet100.
Range Clean Acc. AutoAttack Attack Failure Rate
Swin-T [37] / 86.56 10.64 12.29
Swin-T-AdaNCA [3,5] 87.18 19.52 22.35
[2,6] 87.02 19.02 21.85
[1,7] 86.84 16.20 18.65
C.9 Reason for using ImageNet22K model for RVT
@@
C.10 Visualization of attention maps
@@
C.11 Additional results on same model architecture but with more layers ($\star$ models)
@@
C.12 Additional results on layer similarity structure
@@
Figure 10: Layer similarity structures of different ViTs. Layer set marked in red boxes. AdaNCA is inserted between each pair of stages. $\kappa_{mean}$ is the mean of $\kappa$ of all stages marked out.
C.13 Additional results on category-wise mCE on ImageNet-C
@@
C.14 Additional results on noise sensitivity examination
@@
C.15 Datasets information and license
@@
C.16 Model and code license